OpenStudy (anonymous):

how do I work out pi maths questions say for instance... If a circle has a radius of 3.7, what is the diameter?

OpenStudy (anonymous):

7.4

OpenStudy (shubhamsrg):

pi is just another constant,,like 2,3,10,,etc..its no big deal..

OpenStudy (anonymous):

Yeah but Im not sure how to get the answer can you explain it to me pls

OpenStudy (callisto):

diameter = 2 times radius

OpenStudy (callisto):

Circumference of circle = 2 pi r , where pi is the constant and r is the radius of the circle

OpenStudy (anonymous):

d = 2r where d is the diametre and r is the radius.

OpenStudy (callisto):

Area of the circle = pi r^2 , where pi is the constant and r is the radius of the circle

OpenStudy (anonymous):

Can anyone tell me how to work out the answer on a calculator pls

OpenStudy (anonymous):

jst multiply 3.7*2 and u'll get the answer.

OpenStudy (anonymous):

@Bridgette, @Callisto and @kajolc.2496 gave you the equation. Since the diameter is two times the radius, just multiply the radius by two.

OpenStudy (anonymous):

Thank you very much I have just started learning pi at school and teacher isnt explaining it very well and it is confusing me, im just not getting the pi questions. Is there any other information you could give me to help me understand working out pi questions.

OpenStudy (anonymous):

yaa there are many area of a circle, circumference , diametere,etc

OpenStudy (anonymous):

area = pi*r*r circumference = 2*pi*r diametre = 2r

OpenStudy (anonymous):

Circumference means the length of the outside of a circle. Think of it this way. Make a circle with a piece of string. The length of that string is the length of the circumference. The diameter is the longest line that can fit inside the circle. It always bisects it (cuts it in two). The radius is a straight line from the center point to the outside edge. The diameter is twice the length of the radius. The circumference can be found by taking twice the length of the radius and multiplying by pi. Same thing as the diameter times pi. The area of the circle is the radius squared, and then times pi.

OpenStudy (anonymous):

ok so question.... If a circle has a diameter of 5, what is the radius? r = _____ the answer the wld be 2.5 is this right?

OpenStudy (anonymous):

That is correct.

OpenStudy (anonymous):

If a a circle has radius 7, what is the area? A = _____ wld the answer to this be 153.93804

OpenStudy (anonymous):

Yes, that is correct.

OpenStudy (anonymous):

Thank you so so much you have been a wonderful help I am understanding better now.

OpenStudy (anonymous):

Not a problem at all!